OpenStudy (anonymous):

The value of 6√ lies between which two consecutive integers?

OpenStudy (anonymous):

\[6\sqrt{\text{lies}}\]?

OpenStudy (anonymous):

\[\sqrt{4}\lt\sqrt{6}\lt\sqrt{9}\] \[2\lt\sqrt{6}\lt3\]

OpenStudy (anonymous):

And: \[\{(2\wedge3)\in Z\}\] (2 and 3 are integers)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."